write the slope-intercept form of the equation.
1.) through:(-3,0)and(1,-5)

2.)through(-1,5),perp.to y=1/9x-3

Answer by lwsshak3(11628)

write the slope-intercept form of the equation
1.) through:(-3,0)and(1,-5)
2.)through(-1,5),perp.to y=1/9x-3
Equation of a line: y=mx+b, m=slope , b=y-intercept
1.) through:(-3,0)and(1,-5)
slope,m=∆y/∆x=(-5-0)/(1-(-3))=-5/4
Equation: y=-5x/4+b
solve for b using coordinates of give point given point (-1,5) on the line
-5=-5/4+b
-20/4+5/4=b
b=-15/4
Equation:y=-5x/4-15/4
..
2.)through(-1,5),perp.to y=1/9x-3
y=1/9x-3
slope,m=1/9
slope of perpendicular line=-9 (negative reciprocal)
equation: y=-9x+b
solve for b using coordinates of given point (-1,5) on the line
5=-9*-1+b
b=-4
equation:
y=-9x-4